# Refining Noise Mitigation in NISQ Hardware Through Qubit Error Probability

### v1

@@ Abstract
A key characteristic of the Noisy Intermediate-Scale Quantum (NISQ) era is the accurate assessment and consideration of errors through error mitigation techniques, which play a fundamental role in achieving quantum utility, i.e., quantum results that are at the state of the art of classical calculations. In this paper, we examine the primary sources of errors in contemporary (IBM) quantum computers through a new important metric, the qubit error probability (QEP), which estimates the probability of a qubit to suffer an error. We show that QEP can be used to improve one of the most important error mitigation techniques, the zero noise extrapolation (ZNE). Our method, named zero error probability extrapolation (ZEPE), is based on calibration parameters which enables a good scalability in terms of number of qubits and circuit depth. Besides, we prove that ZEPE performs better than the standard ZNE, specially for a mid-size depth ranges.
  • March 2025
1 Introduction
@@
Quantum computing, defined as the ability to harness real quantum states to perform complex calculations, has recently transitioned from a theoretical possibility to a practical reality [1]. Feynman’s proposition [2] of using quantum systems to simulate quantum mechanics is no longer a conceptual vision but a tangible achievement. Over the past two decades, the capabilities of advanced quantum computers have seen remarkable progress. In principle, ideal quantum computers are expected to execute calculations beyond the reach of classical computers with exceptional precision. However, achieving this accuracy requires devices with thousands of qubits to enable effective quantum error correction [3, 4, 5]. Unfortunately, this goal remains distant. In the current noisy intermediate-scale quantum (NISQ) era, scientific and technological efforts are focused on evaluating, controlling, and mitigating physical errors [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
@@
NISQ devices typically consist of a few dozen qubits and are expected to scale to a few hundred in the near future. A defining characteristic of these devices, as their name implies, is their inherent noisiness. Currently, the most advanced quantum computers are based on transmon qubits [23]. However, these systems face five significant limitations. First, the stability of a qubit state in these devices is typically of the order of hundreds of microseconds, thus restricting the duration of feasible computations. Second, the operations (gates) that act on the qubits are noisy, leading to a progressive loss of accuracy during calculations. Third, the process of measuring the qubit state is prone to errors. Fourth, there is a possibility that an operation on one qubit affects another, the crosstalk. Finally, the number of available physical qubits remains limited, imposing further constraints on computational capability.
@@
Normally, when we want to estimate the error of any quantum calculation, we compare it with the simulated one. Nevertheless, quantum computers in the NISQ era are able to perform state-of-the-art classical computations [24, 25] and it is expected that they may overcome them in the near future. For this reason, it is preferable to avoid relying on classical computations to determine the error in quantum calculations. Therefore, we need to find metrics to measure errors that avoid the use of classical simulations. Sometimes we can use specific Clifford points [24] where the exact analytical result is known and then extrapolate the results to the non-trivial points. Nevertheless, there are many cases in which we may not be able to compare this with any analytical result. In a previous paper [26] we proposed a pre-processing tool (the Tool for Error Description in Quantum Circuits, TED-qc), which allows us to calculate the probability of an error to occur in a quantum circuit provided the hardware and its calibration, without any actual evaluation of the circuit. This tool has proven to be a good estimator of the error of a quantum circuit [27, 28, 29, 30]. However, when a calculation involves many qubits, its usefulness may be limited, as errors in the quantum circuit do not necessarily impact the final result in all cases. In this paper, we introduce a more precise metric for assessing the actual error in a quantum computation. Rather than considering only the total error probability, we focus on the individual qubit error probability (QEP), which provides a more refined measure of the error’s impact on the calculation.
@@
Several error mitigation strategies [31] have been developed to address different sources of errors, including Probabilistic Error Cancellation (PEC), which leverages classical post-processing to counteract noise by applying carefully designed inverse transformations, and Dynamical Decoupling, which use precisely timed control pulses to suppress decoherence and extend qubit coherence times. Additionally, measurement error mitigation techniques correct readout errors by characterizing and inverting the noise model affecting quantum measurements.
Among the error mitigation techniques, one of the most widely used is the Zero-Noise Extrapolation (ZNE) [32, 33, 34], which systematically amplifies noise in a controlled manner, executes the quantum circuit under varying noise regimes, and extrapolates the computed results to approximate an idealized zero-noise limit. This approach relies on the assumption that the quantum system’s response to noise follows a predictable mathematical trend, enabling the reconstruction of an estimated noiseless expectation value. The fundamental principle of ZNE is to employ numerical extrapolation methods —such as linear, polynomial, or exponential regression— to model the dependency of the circuit’s output on noise strength and infer the limit where noise approaches zero. By artificially scaling noise levels, typically through pulse stretching, gate repetition, or hardware-level noise injection, ZNE enables the extraction of noise-free estimates without requiring full quantum error correction. This technique has demonstrated effectiveness in improving the fidelity of quantum computations in domains such as quantum chemistry simulations, variational quantum eigensolvers, and other near-term quantum algorithms where mitigating errors is essential [24, 35, 36, 37]. Importantly, while the computational cost of techniques such as PEC increases exponentially with the number of qubits, ZNE remains independent of qubit count, typically requiring only a error amplification factor of 3-5 in additional quantum computational resources. This scalability makes it one of the most applied quantum error mitigation techniques.
@@
Our benchmark circuit to test both the QEP and its effect in ZNE is the Trotterized time evolution of a two-dimensional transverse-field Ising model. The Ising model is widely studied across various domains of physics and has been extended in innovative ways in recent simulations investigating quantum many-body phenomena, including time crystals, quantum scars, and Majorana edge modes[38, 39, 40, 41].
Specifically, we examine the time dynamics of the Hamiltonian:
@@
where $H_{ZZ}=-J\sum_{\left<i,j\right>}Z_{i}Z_{j}$ and $H_{X}=h\sum_{i}X_{i}$ denote the two separate terms of the Hamiltonian, which are independently exponentiated in the Trotter decomposition. The evolution time $T$ is discretized into $T/\delta t$ Trotter steps, and $R_{Z_{i}Z_{j}}(\theta_{j})$ and $R_{X_{i}}(\theta_{h})$ are ZZ and X rotation gates, respectively. Each Trotter step sequentially applies the two evolution operators defined above, and the total circuit depth increases proportionally with the number of Trotter steps. This flexibility allows for controlled simulation accuracy by increasing the number of steps.
We have chosen $\delta_{t}=1/4$, $J=\pi$, and $h=0$ to position the system at a Clifford point, which enables an efficient comparison between our simulations and the exact analytical solution. The chosen expectation value is the total magnetization $M=\sum_{i}\langle Z_{i}\rangle$.
The paper is organized as follows. In section 2, we introduce the new qubit error probability and compare it with other metrics. In section 3, we introduce a new extrapolation technique, the zero error probability extrapolation (ZEPE) method, to improve standard ZNE. In section 4, we compare the results obtained using our new tool against prior methods. Finally, in section 5, we summarize our conclusions.
2 Qubit error probability
As discussed in the previous section, error estimation is crucial in the NISQ era of quantum computing. In [26], we calculated the total error probability and, following the same scheme, we will calculate the qubit error probability (QEP). As we know, errors arise primarily from four sources:
  • The error induced by the instability of the qubits. In order to have non-trivial states we have to excite our physical qubits. The probability of finding the qubit in the excited state decays exponentially with time [42, 43], so the larger the quantum circuit the less likely is to find the qubit in the expected state. There are two possible decaying mechanisms, the decay of an excited state to the ground state, i.e. the probability of a state $|1\rangle$ to decay to a $|0\rangle$; and the change of the phase of an excited state, for example passing from the state $1/\sqrt{2}(|0\rangle+|1\rangle)$ to the state $1/\sqrt{2}(|0\rangle-|1\rangle)$.
@@ 2 Qubit error probability
  • The crosstalk between different qubits: when we act on a qubit we may act unintentionally on another. This type of error is very difficult to track, but luckily the latest IBM processor, Heron, have almost zero crosstalk [44, 45, 46]. All the results shown in this paper have been calculated on IBM Quantum Heron processor on the $ibm\_torino$ quantum system.
@@
The first term is the measurement error for the j-th qubit, denoted as $P_{j}^{\mathrm{meas}}$, and the second and third terms are the errors arising from qubit instability, represented by $P_{j}^{\tau_{1}}$ and $P_{j}^{\tau_{2}}$. Finally, the error due to all gates applied/affecting on the j-th qubit are considered on the last term, $P_{ij}^{\mathrm{gate}}$. While the gate and measurement errors are directly obtained through the IBM’S API, to account for errors caused by instability qubits $P_{j}^{\tau_{i}}$ we consider an exponential decay of the form [47]
@@
where $i=1,2$ are the two instability mechanism explained before, $t_{j}$ is the total circuit time of j-th qubit and $\tau_{i,j}$ are the relaxation and dephasing times of the j-th qubit. The evaluation of equation 5 requires the knowledge of the time it takes for any qubit between the initialization of the circuit and the measurement $t_{j}$. This is done by adding up how long it takes for the j-th qubit to perform any single gate. The situation becomes more complex when dealing with two-qubit gates. The process of executing a two-qubit gate necessitates the completion of all preceding gate operations on both qubits before the joint operation can be executed. Consequently, after incorporating the time needed for the two-qubit gate into the total time calculation, we are faced with the task of comparing the cumulative times of the two qubits involved. The crucial point here is that we don’t simply add these times together. Instead, we analyze the total times for each qubit and then adopt the longer of the two as the updated qubit time. This approach ensures that we are accounting for the maximum time it could take, thus capturing the full potential for instability-induced error.
@@
These errors depend on the specific quantum hardware, qubits and connections used. To address this, we developed TED-qc (Tool for Error Description in Quantum Circuits), an open-source pre-processing tool available on GitLab (https://gitlab.com/qjornet/ted-qc/). Currenlty written for Qiskit language, it analyzes quantum circuits, extracts error probabilities from IBM’s API and computes total error probabilities. The code can be easily extended to any quantum computer using its calibration parameters.
@@
Figure 1: Schematic representation of the procedure to get the mean Qubit Error Probability. a) 12 phyiscal qubit mapping on $ibm\_torino$ quantum system for the 12 site Ising Hamiltonian (equations 1 and 2). b) to e) Computed Qubit error probability (QEP) of each qubit in a circuit with $1$, $5$, $10$, and $15$ Trotter steps, respectively. Dashed red and green lines represent the mean and standard deviation of the QEP values, respectively.
@@
Figure 1 shows an example of how QEP values are obtained in our tool. We have run a $12$-qubit trotterization of Ising Hamiltonian expressed in equations 1 and 2 for different numbers of Trotter steps on $ibm\_torino$. In the left panel a), we can see the 12-qubit chain chosen by the Qiskit transpile function required to map the logical qubits into the physical quantum hardware. In the right panels, b) to e), we can see the error probability of each qubit in the circuit for $1$, $5$, $10$, and $15$ Trotter steps, respectively. We can see how the mean and variance of the QEP increase as the number of Trotter steps are increased. It is also clear from the image that different qubits have different QEP; however, since the variance is one order of magnitude smaller in all the cases, we take the average value as the extrapolation variable for the error mitigation strategy.
3 Using Qubit Error Probability in the Zero Noise Extrapolation, the ZEPE method
Once we have defined the QEP, we can use it as the extrapolation variable in the ZNE method. We refer to this new ZNE extrapolation method as Zero Error Probability Extrapolation (ZEPE). The idea is the same as in ZNE: we execute multiple circuits with the same expectation value as the original one but ensuring that the replicas will have larger error probabilities each time. To do so, we insert pairs of two-qubit gates on all connected qubits in the circuit. We know that a pair of two-qubit gate is just a unitary operation, so the expectation value remains the same as the original one. The addition of this set of gates defines a factor of one. We can add as many factors as we want to further amplify the noise, and hence the error probability (see figure 2). As we pointed out before, we use the mean QEP to measure the ”error”, and then we extrapolate the mean QEP to zero in order to correct our result.

### v2

# A Useful Metric for the NISQ Era: Qubit Error Probability and Its Role in Zero Noise Extrapolation

@@ Abstract
Accurate assessment and management of errors is indispensable for extracting useful results from noisy intermediate-scale quantum (NISQ) devices. In this work, we propose the qubit error probability (QEP), a device specific metric that combines relaxation, dephasing, gate, and measurement contributions into a single per qubit figure of merit computable before execution. Leveraging QEP as the control variable, we revisit zero noise extrapolation (ZNE) by adding pairs of controlled native two-qubit gates on all connected qubit pairs to generate circuits with successively larger mean QEP; the zero error limit is then approximated by a linear regression of the measured observable against those values.
Benchmarking on IBM Quantum Heron processors, we apply QEP guided ZNE to first order Trotterized simulations of the two dimensional transverse field Ising model, chosen as a representative interacting many body system, involving up to 68 qubits and 15 Trotter steps. In regimes where the raw circuits exhibit a finite mean QEP, the method suppresses observable errors beyond those attainable with circuit depth scaled ZNE, while requiring only three noise scaled evaluations and no additional classical post processing. These results demonstrate that QEP serves as a transparent and efficient error metric, and that its integration into ZNE provides a practical route to reliability gains on current superconducting hardware, without the resource costs associated with full quantum error correction.
preprint: AIP/123-QED
I Introduction
@@
Quantum computing, defined as the ability to harness real quantum states to perform complex calculations, has recently transitioned from a theoretical possibility to a practical reality Nielsen and Chuang (2011). Feynman’s proposition Feynman (1982) of using quantum systems to simulate quantum mechanics is no longer a conceptual vision but a tangible achievement. Over the past two decades, the capabilities of advanced quantum computers have seen remarkable progress. In principle, ideal quantum computers are expected to execute calculations beyond the reach of classical computers with exceptional precision. However, achieving this accuracy requires devices with thousands of qubits to enable effective quantum error correction Lidar and Brun (2013); Terhal (2015); Wendin (2017). Unfortunately, this goal remains distant. In the current noisy intermediate-scale quantum (NISQ) era, scientific and technological efforts are focused on evaluating, controlling, and mitigating physical errors Bharti et al. (2022); Leymann and Barzen (2020); Porter and Joseph (2022); Kandala et al. (2017); Aspuru-Guzik et al. (2005); Cerezo et al. (2021); Arute et al. (2019); Preskill (2018); Xiao, Freericks, and Kemper (2021); Dalzell et al. (2020); Georgopoulos, Emary, and Zuliani (2021); Patel et al. (2020); Nation et al. (2021); Weidenfeller et al. (2022); Setiawan et al. (2021); Wu et al. (2021); Headley et al. (2022).
@@
NISQ devices typically consist of a few dozen qubits and are expected to scale to a few hundred in the near future. A defining characteristic of these devices, as their name implies, is their inherent noisiness. Currently, some of the most advanced quantum computers are based on transmon qubits Koch et al. (2007); Kim et al. (2023); AI and Collaborators (2024). However, these systems face five significant limitations. First, the stability of a qubit state in these devices is typically of the order of hundreds of microseconds, thus restricting the duration of feasible computations. Second, the operations (gates) that act on the qubits are noisy, leading to a progressive loss of accuracy during calculations. Third, the process of measuring the qubit state is prone to errors. Fourth, there is a possibility that an operation on one qubit affects another, the crosstalk. Finally, the number of available physical qubits remains limited, imposing further constraints on computational capability.
@@
The error estimation in a quantum calculation, usually requires comparing the result with the corresponding from a classical calculation. Nevertheless, quantum computers in the NISQ era are able to perform state-of-the-art classical computations Kim et al. (2023); AI and Collaborators (2024) and it is expected that they may overcome them in the near future. For this reason, it is preferable to avoid relying on classical computations to determine the error in quantum calculations. Therefore, we need to find metrics to measure errors that avoid the use of classical simulations. Sometimes we can use specific Clifford points Kim et al. (2023) where the exact analytical result is known and then extrapolate the results to the non-trivial points. However, there are many situations in which it is not possible to compare this with any analytical result. In a previous work Aseguinolaza et al. (2024) we proposed a pre-processing tool (the Tool for Error Description in Quantum Circuits, TED-qc), which computes the probability of an error to occur in a quantum circuit provided the hardware and its calibration, without any actual evaluation of the circuit. This tool has proven to be a good estimator of the error of a quantum circuit Brody and Sykes (2025); Gardeazabal-Gutierrez, Terres-Escudero, and Bringas (2024); Antero et al. (2025); de Keijzer et al. (2025). However, when a calculation involves many qubits, its usefulness may be limited, as errors in the quantum circuit do not necessarily impact the final result in all cases. In this paper, we introduce a more precise metric for assessing the actual error in a quantum computation. Rather than considering only the total error probability, we focus on the individual qubit error probability (QEP), which provides a more refined measure of the error’s impact on the calculation.
@@
Several error mitigation strategies Cai et al. (2023) have been developed to address different sources of error, including Probabilistic Error Cancellation (PEC) Kim et al. (2023), which leverages classical post-processing to counteract noise by applying carefully designed inverse transformations, and Dynamical Decoupling, which use precisely timed control pulses to suppress decoherence and extend qubit coherence times. Additionally, measurement error mitigation techniques correct readout errors by characterizing and inverting the noise model affecting quantum measurements.
Among the error mitigation techniques, one of the most widely used is the Zero-Noise Extrapolation (ZNE) Li and Benjamin (2017); Temme, Bravyi, and Gambetta (2017); Giurgica-Tiron et al. (2020), which systematically amplifies noise in a controlled manner, executes the quantum circuit under varying noise regimes, and extrapolates the computed results to approximate an idealized zero-noise limit. This approach relies on the assumption that the quantum system’s response to noise follows a predictable mathematical trend, enabling the reconstruction of an estimated noiseless expectation value. The fundamental principle of ZNE is to employ numerical extrapolation methods —such as linear, polynomial, or exponential regression— to model the dependency of the circuit’s output on noise strength and infer the limit where noise approaches zero. By artificially scaling noise levels, typically through pulse stretching, gate repetition, or hardware-level noise injection, ZNE enables the extraction of noise-free estimates without requiring full quantum error correction. This technique has demonstrated effectiveness in improving the fidelity of quantum computations in domains such as quantum chemistry simulations, variational quantum eigensolvers, and other near-term quantum algorithms where mitigating errors is essential Kim et al. (2023); Uvarov et al. (2024); Garmon, Pooser, and Dumitrescu (2020); Keen et al. (2020); Cai (2021a); He et al. (2020). Importantly, while the computational cost of techniques such as PEC or Richardson extrapolation Krebsbach, Trauzettel, and Calzona (2022); Russo and Mari (2024) increases exponentially with the number of qubits, ZNE remains independent of qubit count, typically requiring only a error amplification factor of 3-5 in additional quantum computational resources. This scalability makes it one of the most applied quantum error mitigation techniques.
@@
Our benchmark circuit to test both the QEP and its effect in ZNE is the Trotterized time evolution of a two-dimensional transverse-field Ising model. The Ising model is widely studied across various domains of physics and has been extended in innovative ways in recent simulations investigating quantum many-body phenomena, including time crystals, quantum scars, and Majorana edge modesCervera-Lierta (2018); Mi et al. (2022); Frey and Rachel (2022); Chen et al. (2022).
Specifically, we examine the time evolution of the Hamiltonian:
@@
where $H_{ZZ}=-J\sum_{\left<i,j\right>}Z_{i}Z_{j}$ and $H_{X}=h\sum_{i}X_{i}$ denote the two separate terms of the Hamiltonian, which are independently exponentiated in the Trotter decomposition. The evolution time $T$ is discretized into $T/\delta t$ Trotter steps, and $R_{Z_{i}Z_{j}}(\theta_{j})$ and $R_{X_{i}}(\theta_{h})$ are the ZZ and X rotation gates, respectively. Each Trotter step sequentially applies the two evolution operators defined above, and the total circuit depth increases proportionally with the number of Trotter steps. This flexibility allows for controlled simulation accuracy by increasing the number of steps.
Following the lines of Kim Y. et al.Kim et al. (2023) we study the the total magnetization $M=\sum_{i}\langle Z_{i}\rangle$ for two different setup configurations. In the first one, we fix $\delta_{t}=1/4$, $J=\pi$, and $h=0$ to position the system at a Clifford point, allowing for an efficient comparison between our simulations and the exact analytical solution. In this case, we perform calculations up to 68 qubits and vary the number of Trotter steps. In the second one, we fixed the number of Trotter steps to $15$, $\delta_{t}=1/4$, $J=\pi$ for different values of $h$. As we are not at a Clifford point we have done the calculation up to $32$ qubits in order to be able to simulate the theoretical values in a classical computer.
The paper is organized as follows. In section II, we introduce the qubit error probability (QEP), and compare it with other metrics. In section III, we use this QEP to track the error in order to improve standard ZNE. In section IV, we compare the results obtained using our new tool against prior methods. Finally, in section V, we summarize our conclusions.
II Qubit error probability
As discussed in the previous section, error estimation is crucial in the NISQ era of quantum computing. In Aseginolaza U. et al.Aseguinolaza et al. (2024), we calculated the total error probability and, following the same scheme, we will calculate the qubit error probability (QEP). As we know, errors arise primarily from four sources:
  • The error induced by the instability of the qubits. In order to have non-trivial states we have to excite our physical qubits. The probability of finding the qubit in the excited state decays exponentially with time Rigetti et al. (2012); Place et al. (2021), so the larger the quantum circuit the less likely is to find the qubit in the expected state. There are two possible decaying mechanisms, the decay of an excited state to the ground state, i.e. the probability of a state $|1\rangle$ to decay to a $|0\rangle$; and the change of the phase of an excited state, for example passing from the state $1/\sqrt{2}(|0\rangle+|1\rangle)$ to the state $1/\sqrt{2}(|0\rangle-|1\rangle)$.
@@ II Qubit error probability
  • The crosstalk between different qubits: when we act on a qubit we may act unintentionally on another. This type of error is very difficult to track, but luckily the latest IBM processor, Heron, have almost zero crosstalk Sheldon et al. (2016); Dial (2022); Chow, Dial, and Gambetta (2021). All the results shown in this paper have been calculated on IBM Quantum Heron processor on the $ibm\_torino$ quantum system.
@@
The first term is the measurement error for the j-th qubit, denoted as $P_{j}^{\mathrm{meas}}$, and the second and third terms are the errors arising from qubit instability, represented by $P_{j}^{\tau_{1}}$ and $P_{j}^{\tau_{2}}$. Finally, the error due to all gates applied/affecting on the j-th qubit are considered on the last term, $P_{ij}^{\mathrm{gate}}$. While the gate and measurement errors are directly obtained through the IBM’s API, to account for errors caused by instability qubits $P_{j}^{\tau_{i}}$ we consider an exponential decay of the form Aleksandrowicz et al. (2019)
@@
where $i=1,2$ stand for the two instability mechanism, relaxation and dephasing explained before, $\tau_{i,j}$ are the related times for the j-th qubit, and $t_{j}$ is the total circuit time of j-th qubit. The evaluation of equation 5 requires the knowledge of the time it takes for any qubit between the initialization of the circuit and the measurement $t_{j}$. This is done by adding up how long it takes for the j-th qubit to perform any single gate. The situation becomes more complex when dealing with two-qubit gates. The process of executing a two-qubit gate necessitates the completion of all preceding gate operations on both qubits before the joint operation can be executed. Consequently, after incorporating the time needed for the two-qubit gate into the total time calculation, we are faced with the task of comparing the cumulative times of the two qubits involved. The crucial point here is that we don’t simply add these times together. Instead, we analyze the total times for each qubit and then adopt the longer of the two as the updated qubit time. This approach ensures that we are accounting for the maximum time it could take, thus capturing the full potential for instability-induced error.
@@
These errors depend on the specific quantum hardware, qubits and connections used. To address this, we developed TED-qc (Tool for Error Description in Quantum Circuits), an open-source pre-processing tool available on GitLab (https://gitlab.com/qjornet/ted-qc/). Currenlty written for Qiskit v1.2.0 language, it analyzes quantum circuits, extracts error probabilities from IBM’s API and computes total error probabilities. The code can be easily extended to any quantum computer using its calibration parameters.
@@
Figure 1: Schematic representation of the procedure to get the mean Qubit Error Probability. a) 12 phyiscal qubit mapping on $ibm\_torino$ quantum system for the 12 site Ising Hamiltonian (equation 1). b) to e) Computed Qubit error probability (QEP) of each qubit in a circuit with $1$, $5$, $10$, and $15$ Trotter steps, respectively. Dashed red and green lines represent the mean ($\mu$) and standard deviation ($\sigma$) of the QEP values, respectively.
@@
Figure 1 shows an example of how QEP values are obtained in our tool. We have run a $12$-qubit trotterization of Ising Hamiltonian expressed in equation 1 with $\delta_{t}=1/4$, $J=\pi$, and $h=0$ for different number of Trotter steps on $ibm\_torino$. In the left panel a), we can see the 12-qubit chain chosen by the Qiskit transpile function required to map the logical qubits into the physical quantum hardware. In the right panels, b) to e), we can see the error probability of each qubit in the circuit for $1$, $5$, $10$, and $15$ Trotter steps, respectively. We can see how the mean ($\mu$) and variance ($\sigma$) of the QEP increase as the number of Trotter steps are increased. It is also clear from the image that different qubits have different QEP; however, since the variance is one order of magnitude smaller in all the cases, we take the average value as the extrapolation variable for the error mitigation strategy.
III Using Qubit Error Probability in the Zero Noise Extrapolation